A merchant has coffee worth $4 a pound that she wishes to mix with 80 pounds of coffee worth $9 a pound to get a mixture that can be sold for $5 a pound. How many pounds of the $4 coffee should be used?
My answer = 320 lbs
Thanks for your help.

Answer by nerdybill(7384)

A merchant has coffee worth $4 a pound that she wishes to mix with 80 pounds of coffee worth $9 a pound to get a mixture that can be sold for $5 a pound. How many pounds of the $4 coffee should be used?
.
You would be correct!
.
Let x = amount (pounds) of $4 coffee
then
4x + 9(80) = 5(x+80)
4x + 720 = 5x+400
720 = x+400
320 pounds = x